# Breaking the Assistant Mold: Modeling Behavioral Variation in LLM Based Procedural Character Generation

### v1

@@ 1 Introduction
Figure 1: Illustration of biases in prior work versus our method Prior work populations jinworldweaver in (a)-Top collapse into uniform agreement with moral statements from the Social Chemistry dataset forbes-etal-2020-social, reflecting a moral bias, while in (a)-Bottom they invariably answer conversational questions from ConvAI2 dinan2019second, reflecting a reaction bias. In contrast, our generated populations (b)-Top and (b)-Bottom exhibit a broader range of stances and responses.
Procedural character generation (PCG) is a recent yet underexplored task that aims to automatically create diverse and believable agents for virtual and narrative environments. Enabled by advances in large language models (LLMs), which unlock new possibilities for character role play wang2023rolellm; zhou2023characterglm; shao2023character, LLMs are an attractive foundation for PCG. Existing approaches—either directly prompting LLMs to produce populations jinworldweaver or adapting scraped persona banks to specific settings ge2024scaling—inherit alignment-induced biases: a positive moral bias, where characters uniformly adopt agreeable stances (Fig. 1 Top), and a helpful assistant bias, where they invariably answer questions (Fig. 1 Bottom). While such tendencies are desirable in instruction-following systems, they suppress dramatic tension and lead to repetitive archetypes. This reflects a broader limitation of maximum-likelihood training and assistant fine-tuning, which bias models toward safe continuations and homogenized voices, as also observed in other simulation tasks kotek2023gender; cheng2023marked; wang2025large.
@@
To address these issues, we introduce PersonaWeaver, which tackles the collapse of behavior by disentangling world-building from behavioral-building (Fig. 2). This separation allows us to explicitly model variation in behavior, providing finer control to probe whether LLMs can realize diverse dispositions. Unlike world-building, which is setting-specific (e.g., a corn farmer does not live in New York City), behavioral traits such as moral stances and interactional styles are universal, making them amenable to external banks that allow systematic specification. For example, we use Moral Foundations Theory graham2013moral to define stances across axes such as care, fairness, and authority, and an interactional bank to capture conversational tendencies such as cooperation, disagreement, or evasion. By sampling and recombining these dimensions, PersonaWeaver creates character profiles that exhibit varied reactions and moral stances (Fig. 2). In sum, our results demonstrate that PersonaWeaver elicits substantially more diverse responses than prior work when behavior is explicitly modeled (Fig. 1 b).
We demonstrate this through a large-scale experimental procedure, instantiating character profiles across fictional and non-fictional settings and testing them on three frontier models: GPT-4o achiam2023gpt, LLaMA 4 meta2024llama4, and Qwen 3 yang2025qwen3. Characters are then probed on two tasks. In a moral belief experiment adapted from Social Chemistry forbes-etal-2020-social, they are presented with normative statements (e.g., , Fig. 1 a) and asked to register agreement or disagreement; PersonaWeaver yields a broader, more balanced distribution of stances than prior work, avoiding collapse to uniform agreement. In an interaction experiment using conversational prompts from ConvAI2 dinan2019second, characters produce open-ended responses; here, PersonaWeaver exhibits more varied behaviors—refusal, deflection, disagreement—while also inducing second-order stylistic diversity in features such as length, punctuation, and emotional tone, resulting in richer and less homogenized profiles.
@@ 2 A Tale of Two Biases in PCG
We are concerned with identifying and mitigating two types of behavioral biases that arise in existing approaches to LLM-based procedural character generation jinworldweaver; ge2024scaling: (1) positive moral bias, where characters overwhelmingly agree on normative statements, and (2) helpful assistant bias, where characters invariably answer questions directly. To this end, we define our study preliminaries in Sec. 2.1, empirically demonstrate the two biases in Sec. 2.2, and introduce PersonaWeaver in Sec. 2.3 and show how it mitigates them.
@@ 2.1 Preliminaries
Prior Work. WorldWeaver jinworldweaver generates characters through direct prompting of the LLM (e.g., Generate $N$ different Character profiles for Setting $T$). PersonaHub ge2024scaling samples personas from an internet-scale bank of scraped profiles and adapts them to the target setting using an LLM.
Simulations. Prior work jinworldweaver; ge2024scaling has evaluated character generation within a single setting, limiting insight into how methods generalize across diverse contexts. Thus, we construct a broader experimental environment by instantiating populations across 10 distinct settings (five fictional, five real-life), sampled from popular movies and television to ensure cultural and geographic diversity. For each setting, we generate 100 characters, yielding 1,000 characters per method. Refer to Appendix E for further details.
Figure 3: Comparison of Moral Positions between prior work WorldWeaver jinworldweaver, PersonaHub ge2024scaling and our method PersonaHub. Refer to Sec 2.2 and 2.3.1 for discussion.
Figure 4: Comparison of Reactions to Questions between prior work WorldWeaver jinworldweaver, PersonaHub ge2024scaling and our method PersonaHub. Refer to Sec 2.2 and 2.3.1 for discussion.
Quantifying Biases. We examine moral bias, where characters uniformly adopt agreeable stances, and reaction bias, where they invariably answer questions directly. To probe moral bias, Per Fig. 1 Top, we present characters with moral norm statements from Social Chemistry forbes-etal-2020-social and measure the distribution of multiple-choice responses. To probe reaction bias, Per Fig. 1 Bottom, we ask characters conversational questions drawn from ConvAI2 dinan2019second ; their open-ended replies are classified by an auxiliary LLM into three categories (refusal, deflection, compliance), yielding distributions over behaviors. Refer to Appendix B for further details.
Models. Our results in the paper are based on GPT-4o achiam2023gpt, a frontier LLM. To ensure generality, we replicate all experiments on LLaMA 4 meta2024llama4 and Qwen 3 yang2025qwen3; in Appendix F.
@@
Figure 5: Comparison of Stylistic Patterns in the generated answers of prior work (WorldWeaver jinworldweaver and PersonaHub ge2024scaling) and our work PersonaWeaver across four Stylistic categories (Filler words, Punctuations, Answer Length, and Sentiment). Refer to Section 2.3.1 for further discussion.
@@ 2.2 Demonstrating Biases Empirically
Discussion. These results show how alignment-driven objectives—maximum likelihood estimation, which favors high-probability continuations, and assistant fine-tuning ouyang2022training, which rewards helpfulness—collapse behavioral diversity. Characters thus inherit a moral bias toward universal agreement and a reaction bias toward always answering, reinforcing predictable, assistant-like voices. In Sec. 2.3, we introduce PersonaWeaver, designed to probe and mitigate these biases.
@@ 2.3 PersonaWeaver: Disentangling World from Behavior Building
Behavioral Module. Unlike world-building, which is setting-dependent, behavioral traits are universal, so we define them through external banks that give us systematic control. For moral traits, we draw inspiration from Moral Foundations Theory graham2013moral, defining eight distinct moral stances that are sampled and attached to character profiles, producing systematic variation in how characters judge social norms (full list in the Appendix C). For interactional traits, we define eight categories spanning a spectrum of conversational behaviors: refusals, deflections, hesitations, direct compliance, volunteering additional details, playful or subversive replies, hostile retorts, and meta-level commentary. Together, these banks allow us to break the “always answer” defaults that dominate prior work. Refer to Appendix D for further details.
@@ 4 Limitations
While PersonaWeaver broadens behavioral diversity in generated characters, our evaluation is constrained to moral and interactional dimensions. Extending this framework to other behavioral axes—such as emotional regulation—remains an open direction and can be easily investigated with our framework. Additionally, our moral evaluation relies on multiple-choice agreement tasks, which provide controlled but coarse-grained signals of moral stance. Such format may underrepresent nuanced or context-dependent reasoning that could emerge in richer, free-form moral dialogue, suggesting a need for more naturalistic evaluations in future work.
Potential Risks Expanding behavioral diversity in generated characters introduces ethical risks, particularly when simulating morally or socially sensitive behaviors. While PersonaWeaver aims to probe alignment biases, increased variation could inadvertently surface undesirable or offensive responses if not properly constrained. Furthermore, explicit control over moral and interactional traits raises concerns about reinforcing or exploiting stereotypes if applied uncritically. Careful curation of behavioral banks and context-aware deployment are essential to ensure that diversity serves creative and research goals without amplifying harm.
@@ Appendix A Related Works
Procedural Character Generation. Research on procedural character generation remains limited, with most prior work focusing on producing characters within a single, predefined environment (e.g., , one game world jinworldweaver; ge2024scaling). WorldWeaver jinworldweaver prompts an LLM to directly generate $N$ characters for a given setting, while PersonaHub ge2024scaling samples personas from large-scale scraped profile datasets and adapts them to the target context using an LLM. In this work, we show that these approaches struggle with two key issues—limited moral diversity and restricted interactional range—and introduce a method designed to explicitly mitigate these biases.
Biases in Simulating Personas. Prior studies have documented a broader lack of behavioral and identity diversity in simulated personas. For instance, Marked Personas cheng2023marked show that language models tend to reproduce social stereotypes, while others highlight gender and identity flattening kotek2023gender; wang2025large and reduced narrative variety in multimodal storytelling lee2024vision. Such homogenization arises from maximum-likelihood training objectives that favor high-probability continuations, as well as alignment fine-tuning that rewards politeness, safety, and helpfulness—thereby suppressing the variability essential to believable social simulation. In our work, we extend this observation to procedural character generation, showing that similar alignment-induced biases constrain moral and interactional diversity in LLM-generated populations.
Character Simulation and Role-Playing. A growing line of research explores how LLMs can function as conversational agents with consistent personality, memory, and long-term coherence. Generative Agents park2023generative simulate memory-driven individuals inhabiting a shared environment, though their backgrounds are largely hand-crafted. In the role-playing domain, works such as RoleLLM wang2023rolellm, CharacterGLM zhou2023characterglm, and Character-LLM shao2023character focus on eliciting and sustaining role-specific behaviors through persona-conditioned dialogue systems. These efforts primarily target the believability and consistency of agent role play. In contrast, our work examines whether off-the-shelf LLMs can leverage their broad world knowledge to simulate behaviorally diverse populations—spanning distinct moral dispositions and interactional tendencies—rather than a few coherent yet homogeneous individuals.
@@
Social Chemistry forbes-etal-2020-social Parents are expected to make sure their kids eat healthy food.
@@
ConvAI2 dinan2019second Sentiment (s1–s5):
@@
Table 1: Moral statements and conversational questions used in our evaluation. Moral norms are drawn from Social Chemistry forbes-etal-2020-social, while conversational prompts are extracted from ConvAI2 dinan2019second. Refer to Appendix B for further discussion.
@@
Table 2: The eight moral positions used in our evaluation, inspired by Moral Foundations Theory graham2013moral. Each position encodes a distinct stance toward compassion, fairness, loyalty, authority, purity, and liberty. Refer to Appendix C for further discussion.
@@
Procedural Content Generation with LLMs. Beyond characters, LLMs have been used for procedural generation of levels, stories, and worlds. For example, Word2World nasir2024word2world generates narratives/world descriptions, Mariogpt sudhakaran2023mariogpt generates levels, and works in text adventure or interactive story generation freiknecht2020procedural; hu2024game show how LLMs can produce dynamic environments. WorldWeaver jinworldweaver also overlaps somewhat, using LLMs to generate world contexts and roles.
@@ Appendix B Dataset Details
Our evaluations draw on two sources of prompts: moral norm statements from Social Chemistry forbes-etal-2020-social and conversational questions from ConvAI2 dinan2019second. Table 1 summarizes the full set used in our experiments.
Social Chemistry. We use a curated set of everyday moral statements from the Social Chemistry dataset forbes-etal-2020-social, which encodes widely held social norms (e.g., “Parents are expected to make sure their kids eat healthy food”). These statements serve as probes for whether generated characters adopt diverse moral stances rather than collapsing into uniform agreement.
ConvAI2. For conversational prompts, we extract candidate utterances from ConvAI2 dinan2019second dialogues by filtering for sentences that end with a question mark. From these, we select two categories: general-purpose questions (q1–q5) about hobbies, preferences, or opinions, and sentiment-oriented probes (s1–s5) about feelings or states. This set allows us to examine whether generated characters vary in their interactional responses (e.g., refusal, deflection, compliance) rather than defaulting to always answering.
@@ Appendix C Moral Positions
To model systematic variation in moral reasoning, we define eight moral positions inspired by Moral Foundations Theory graham2013moral. Each position encodes a distinct configuration of stances toward compassion, fairness, loyalty, authority, purity, and liberty.
@@ Appendix F Additional Results: LLaMA and Qwen
In the main paper, we presented results using GPT-4o achiam2023gpt as the primary model of study. Specifically, we examined (i) the distribution of moral stances elicited by normative statements (Fig. 3), (ii) the distribution of reactions to open-ended conversational questions (Fig. 4), and (iii) second-order stylistic effects such as filler word usage, punctuation, answer length, and sentiment (Fig. 5). These experiments established the two alignment-induced biases of existing methods and demonstrated how PersonaWeaver mitigates them by inducing broader behavioral and stylistic diversity.
@@
Figure 6: Comparison of Moral and Reaction Biases across two additional models. Top row: LLaMA. Bottom row: Qwen. Each pair contrasts prior work (WorldWeaver jinworldweaver, PersonaHub ge2024scaling) with our approach PersonaWeaver. Refer to Section 2.2 and 2.3.1 for discussion.
LLaMA 4 meta2024llama4
@@
Qwen 3 yang2025qwen3
Figure 7: Comparison of Stylistic Patterns in the generated answers of prior work (WorldWeaver jinworldweaver and PersonaHub ge2024scaling) and our work PersonaWeaver across four stylistic categories (Filler words, Punctuations, Answer Length, and Sentiment). Results are shown for LLaMA 4 meta2024llama4 (top row) and Qwen 3 yang2025qwen3 (bottom row). Refer to Section 2.3.1 for further discussion.
Here, we replicate these analyses using two additional frontier models: LLaMA 4 meta2024llama4 and Qwen 3 yang2025qwen3. Fig. 6 reports the distributions of moral stances and conversational reactions, while Fig. 7 extends the stylistic comparison across filler words, punctuation, length, and sentiment. Across both models, we observe patterns consistent with those documented for GPT-4o achiam2023gpt: prior methods (WorldWeaver jinworldweaver, PersonaHub ge2024scaling) collapse into agreement and compliance, while PersonaWeaver broadens the space of moral positions, elicits more varied reactions (e.g., refusals, deflections, playful responses), and induces second-order stylistic diversity.

### v2

@@ 1 Introduction
Figure 1: Illustration of biases in prior work versus our method Prior work populations Jin et al. (2024) in (a)-Top collapse into uniform agreement with moral statements from the Social Chemistry dataset Forbes et al. (2020), reflecting a moral bias, while in (a)-Bottom they invariably answer conversational questions from ConvAI2 Dinan et al. (2019), reflecting a reaction bias. In contrast, our generated populations (b)-Top and (b)-Bottom exhibit a broader range of stances and responses.
Procedural character generation (PCG) is a recent yet underexplored task that aims to automatically create diverse and believable agents for virtual and narrative environments. Enabled by advances in large language models (LLMs), which unlock new possibilities for character role play Wang et al. (2023); Zhou et al. (2023); Shao et al. (2023), LLMs are an attractive foundation for PCG. Existing approaches—either directly prompting LLMs to produce populations Jin et al. (2024) or adapting scraped persona banks to specific settings Ge et al. (2024)—inherit alignment-induced biases: a positive moral bias, where characters uniformly adopt agreeable stances (Fig. 1 Top), and a helpful assistant bias, where they invariably answer questions (Fig. 1 Bottom). While such tendencies are desirable in instruction-following systems, they suppress dramatic tension and lead to repetitive archetypes. This reflects a broader limitation of maximum-likelihood training and assistant fine-tuning, which bias models toward safe continuations and homogenized voices, as also observed in other simulation tasks Kotek et al. (2023); Cheng et al. (2023); Wang et al. (2025).
@@
To address these issues, we introduce PersonaWeaver, which tackles the collapse of behavior by disentangling world-building from behavioral-building (Fig. 2). This separation allows us to explicitly model variation in behavior, providing finer control to probe whether LLMs can realize diverse dispositions. Unlike world-building, which is setting-specific (e.g., a corn farmer does not live in New York City), behavioral traits such as moral stances and interactional styles are universal, making them amenable to external banks that allow systematic specification. For example, we use Moral Foundations Theory Graham et al. (2013) to define stances across axes such as care, fairness, and authority, and an interactional bank to capture conversational tendencies such as cooperation, disagreement, or evasion. By sampling and recombining these dimensions, PersonaWeaver creates character profiles that exhibit varied reactions and moral stances (Fig. 2). In sum, our results demonstrate that PersonaWeaver elicits substantially more diverse responses than prior work when behavior is explicitly modeled (Fig. 1 b).
We demonstrate this through a large-scale experimental procedure, instantiating character profiles across fictional and non-fictional settings and testing them on three frontier models: GPT-4o Achiam et al. (2023), LLaMA 4 AI (2024), and Qwen 3 Yang et al. (2025). Characters are then probed on two tasks. In a moral belief experiment adapted from Social Chemistry Forbes et al. (2020), they are presented with normative statements (e.g., , Fig. 1 a) and asked to register agreement or disagreement; PersonaWeaver yields a broader, more balanced distribution of stances than prior work, avoiding collapse to uniform agreement. In an interaction experiment using conversational prompts from ConvAI2 Dinan et al. (2019), characters produce open-ended responses; here, PersonaWeaver exhibits more varied behaviors—refusal, deflection, disagreement—while also inducing second-order stylistic diversity in features such as length, punctuation, and emotional tone, resulting in richer and less homogenized profiles.
@@ 2 A Tale of Two Biases in PCG
We are concerned with identifying and mitigating two types of behavioral biases that arise in existing approaches to LLM-based procedural character generation Jin et al. (2024); Ge et al. (2024): (1) positive moral bias, where characters overwhelmingly agree on normative statements, and (2) helpful assistant bias, where characters invariably answer questions directly. To this end, we define our study preliminaries in Sec. 2.1, empirically demonstrate the two biases in Sec. 2.2, and introduce PersonaWeaver in Sec. 2.3 and show how it mitigates them.
@@ 2.1 Preliminaries
Prior Work. WorldWeaver Jin et al. (2024) generates characters through direct prompting of the LLM (e.g., Generate $N$ different Character profiles for Setting $T$). PersonaHub Ge et al. (2024) samples personas from an internet-scale bank of scraped profiles and adapts them to the target setting using an LLM.
Simulations. Prior work Jin et al. (2024); Ge et al. (2024) has evaluated character generation within a single setting, limiting insight into how methods generalize across diverse contexts. Thus, we construct a broader experimental environment by instantiating populations across 10 distinct settings (five fictional, five real-life), sampled from popular movies and television to ensure cultural and geographic diversity. For each setting, we generate 100 characters, yielding 1,000 characters per method. Refer to Appendix E for further details.
Figure 3: Comparison of Moral Positions between prior work WorldWeaver Jin et al. (2024), PersonaHub Ge et al. (2024) and our method PersonaHub. Refer to Sec 2.2 and 2.3.1 for discussion.
Figure 4: Comparison of Reactions to Questions between prior work WorldWeaver Jin et al. (2024), PersonaHub Ge et al. (2024) and our method PersonaHub. Refer to Sec 2.2 and 2.3.1 for discussion.
Quantifying Biases. We examine moral bias, where characters uniformly adopt agreeable stances, and reaction bias, where they invariably answer questions directly. To probe moral bias, Per Fig. 1 Top, we present characters with moral norm statements from Social Chemistry Forbes et al. (2020) and measure the distribution of multiple-choice responses. To probe reaction bias, Per Fig. 1 Bottom, we ask characters conversational questions drawn from ConvAI2 Dinan et al. (2019) ; their open-ended replies are classified by an auxiliary LLM into three categories (refusal, deflection, compliance), yielding distributions over behaviors. Refer to Appendix B for further details.
Models. Our results in the paper are based on GPT-4o Achiam et al. (2023), a frontier LLM. To ensure generality, we replicate all experiments on LLaMA 4 AI (2024) and Qwen 3 Yang et al. (2025); in Appendix F.
@@
Figure 5: Comparison of Stylistic Patterns in the generated answers of prior work (WorldWeaver Jin et al. (2024) and PersonaHub Ge et al. (2024)) and our work PersonaWeaver across four Stylistic categories (Filler words, Punctuations, Answer Length, and Sentiment). Refer to Section 2.3.1 for further discussion.
@@ 2.2 Demonstrating Biases Empirically
Discussion. These results show how alignment-driven objectives—maximum likelihood estimation, which favors high-probability continuations, and assistant fine-tuning Ouyang et al. (2022), which rewards helpfulness—collapse behavioral diversity. Characters thus inherit a moral bias toward universal agreement and a reaction bias toward always answering, reinforcing predictable, assistant-like voices. In Sec. 2.3, we introduce PersonaWeaver, designed to probe and mitigate these biases.
@@ 2.3 PersonaWeaver: Disentangling World from Behavior Building
Behavioral Module. Unlike world-building, which is setting-dependent, behavioral traits are universal, so we define them through external banks that give us systematic control. For moral traits, we draw inspiration from Moral Foundations Theory Graham et al. (2013), defining eight distinct moral stances that are sampled and attached to character profiles, producing systematic variation in how characters judge social norms (full list in the Appendix C). For interactional traits, we define eight categories spanning a spectrum of conversational behaviors: refusals, deflections, hesitations, direct compliance, volunteering additional details, playful or subversive replies, hostile retorts, and meta-level commentary. Together, these banks allow us to break the “always answer” defaults that dominate prior work. Refer to Appendix D for further details.
@@ 4 Limitations
Out evaluation is limited in two major ways: 1) it only examines two behavioral dimensions: moral stances and characters’ interactions. It overlooks dimensions like emotional regulation. Therefore, future work can benefit from expanding our evaluation setup 2) While our coarse grained moral stance evaluation enables a streamlined study, it overlooks more nuanced moral reasoning that can’t be simply coarsely categorized.
Potential Risks Efforts to systemically expand behavioral diversity in character generation can push models to generate characters that replicate offensive or unsafe behaviors. Therefore, careful curation of behavioral banks is essential to ensure that increasing diversity in character generation serves creative and research goals without amplifying harm.
@@ Appendix A Related Works
Procedural Character Generation. Research on procedural character generation remains limited, with most prior work focusing on producing characters within a single, predefined environment (e.g., , one game world Jin et al. (2024); Ge et al. (2024)). WorldWeaver Jin et al. (2024) prompts an LLM to directly generate $N$ characters for a given setting, while PersonaHub Ge et al. (2024) samples personas from large-scale scraped profile datasets and adapts them to the target context using an LLM. In this work, we show that these approaches struggle with two key issues—limited moral diversity and restricted interactional range—and introduce a method designed to explicitly mitigate these biases.
Biases in Simulating Personas. Prior studies have documented a broader lack of behavioral and identity diversity in simulated personas. For instance, Marked Personas Cheng et al. (2023) show that language models tend to reproduce social stereotypes, while others highlight gender and identity flattening Kotek et al. (2023); Wang et al. (2025) and reduced narrative variety in multimodal storytelling Lee and Jeon (2024). Such homogenization arises from maximum-likelihood training objectives that favor high-probability continuations, as well as alignment fine-tuning that rewards politeness, safety, and helpfulness—thereby suppressing the variability essential to believable social simulation. In our work, we extend this observation to procedural character generation, showing that similar alignment-induced biases constrain moral and interactional diversity in LLM-generated populations.
Character Simulation and Role-Playing. A growing line of research explores how LLMs can function as conversational agents with consistent personality, memory, and long-term coherence. Generative Agents Park et al. (2023) simulate memory-driven individuals inhabiting a shared environment, though their backgrounds are largely hand-crafted. In the role-playing domain, works such as RoleLLM Wang et al. (2023), CharacterGLM Zhou et al. (2023), and Character-LLM Shao et al. (2023) focus on eliciting and sustaining role-specific behaviors through persona-conditioned dialogue systems. These efforts primarily target the believability and consistency of agent role play. In contrast, our work examines whether off-the-shelf LLMs can leverage their broad world knowledge to simulate behaviorally diverse populations—spanning distinct moral dispositions and interactional tendencies—rather than a few coherent yet homogeneous individuals.
@@
Social Chemistry Forbes et al. (2020) Parents are expected to make sure their kids eat healthy food.
@@
ConvAI2 Dinan et al. (2019) Sentiment (s1–s5):
@@
Table 1: Moral statements and conversational questions used in our evaluation. Moral norms are drawn from Social Chemistry Forbes et al. (2020), while conversational prompts are extracted from ConvAI2 Dinan et al. (2019). Refer to Appendix B for further discussion.
@@
Table 2: The eight moral positions used in our evaluation, inspired by Moral Foundations Theory Graham et al. (2013). Each position encodes a distinct stance toward compassion, fairness, loyalty, authority, purity, and liberty. Refer to Appendix C for further discussion.
@@
Procedural Content Generation with LLMs. Beyond characters, LLMs have been used for procedural generation of levels, stories, and worlds. For example, Word2World Nasir et al. (2024) generates narratives/world descriptions, Mariogpt Sudhakaran et al. (2023) generates levels, and works in text adventure or interactive story generation Freiknecht and Effelsberg (2020); Hu et al. (2024) show how LLMs can produce dynamic environments. WorldWeaver Jin et al. (2024) also overlaps somewhat, using LLMs to generate world contexts and roles.
@@ Appendix B Dataset Details
Our evaluations draw on two sources of prompts: moral norm statements from Social Chemistry Forbes et al. (2020) and conversational questions from ConvAI2 Dinan et al. (2019). Table 1 summarizes the full set used in our experiments.
Social Chemistry. We use a curated set of everyday moral statements from the Social Chemistry dataset Forbes et al. (2020), which encodes widely held social norms (e.g., “Parents are expected to make sure their kids eat healthy food”). These statements serve as probes for whether generated characters adopt diverse moral stances rather than collapsing into uniform agreement.
ConvAI2. For conversational prompts, we extract candidate utterances from ConvAI2 Dinan et al. (2019) dialogues by filtering for sentences that end with a question mark. From these, we select two categories: general-purpose questions (q1–q5) about hobbies, preferences, or opinions, and sentiment-oriented probes (s1–s5) about feelings or states. This set allows us to examine whether generated characters vary in their interactional responses (e.g., refusal, deflection, compliance) rather than defaulting to always answering.
@@ Appendix C Moral Positions
To model systematic variation in moral reasoning, we define eight moral positions inspired by Moral Foundations Theory Graham et al. (2013). Each position encodes a distinct configuration of stances toward compassion, fairness, loyalty, authority, purity, and liberty.
@@ Appendix F Additional Results: LLaMA and Qwen
In the main paper, we presented results using GPT-4o Achiam et al. (2023) as the primary model of study. Specifically, we examined (i) the distribution of moral stances elicited by normative statements (Fig. 3), (ii) the distribution of reactions to open-ended conversational questions (Fig. 4), and (iii) second-order stylistic effects such as filler word usage, punctuation, answer length, and sentiment (Fig. 5). These experiments established the two alignment-induced biases of existing methods and demonstrated how PersonaWeaver mitigates them by inducing broader behavioral and stylistic diversity.
@@
Figure 6: Comparison of Moral and Reaction Biases across two additional models. Top row: LLaMA. Bottom row: Qwen. Each pair contrasts prior work (WorldWeaver Jin et al. (2024), PersonaHub Ge et al. (2024)) with our approach PersonaWeaver. Refer to Section 2.2 and 2.3.1 for discussion.
LLaMA 4 AI (2024)
@@
Qwen 3 Yang et al. (2025)
Figure 7: Comparison of Stylistic Patterns in the generated answers of prior work (WorldWeaver Jin et al. (2024) and PersonaHub Ge et al. (2024)) and our work PersonaWeaver across four stylistic categories (Filler words, Punctuations, Answer Length, and Sentiment). Results are shown for LLaMA 4 AI (2024) (top row) and Qwen 3 Yang et al. (2025) (bottom row). Refer to Section 2.3.1 for further discussion.
Here, we replicate these analyses using two additional frontier models: LLaMA 4 AI (2024) and Qwen 3 Yang et al. (2025). Fig. 6 reports the distributions of moral stances and conversational reactions, while Fig. 7 extends the stylistic comparison across filler words, punctuation, length, and sentiment. Across both models, we observe patterns consistent with those documented for GPT-4o Achiam et al. (2023): prior methods (WorldWeaver Jin et al. (2024), PersonaHub Ge et al. (2024)) collapse into agreement and compliance, while PersonaWeaver broadens the space of moral positions, elicits more varied reactions (e.g., refusals, deflections, playful responses), and induces second-order stylistic diversity.